Fred weighs 187 pounds but is on a diet that makes him lose 1.7 pounds per week. Barney weighs only 93 pounds but is on a diet that makes him gain 0.9 pounds per week. What is the number of weeks it takes for Fred and Barney to weigh exactly the same amount? How many weeks does it take for their total weight to equal 252 pounds?
Answer by stanbon(75887)

Fred weighs 187 pounds but is on a diet that makes him lose 1.7 pounds per week.
F(x) = 187-1.7x
-------
Barney weighs only 93 pounds but is on a diet that makes him gain 0.9 pounds per week.
B(x) = 93+0.9x
-------
What is the number of weeks it takes for Fred and Barney to weigh exactly the same amount?
Solve F(x) = B(x) for "x":
187-1.7x = 93+0.9x
2.6x = 194
x = 74.62
Equal weight during the 74th week,
================
How many weeks does it take for their total weight to equal 252 pounds?
Solve F(x) + B(x) = 252
187-1.7x + 83+0.9x = 252
-0.8x + 270 = 252
-0.8x = -18
x = 22.5 weeks
=============================
Cheers,
Stan H.
================